OpenStudy (anonymous):

Fully factorize -6x^2+12x

OpenStudy (anonymous):

if you take 6x out what do you get ?

OpenStudy (anonymous):

x+12x 1*x+1*12*x 1x(12) ? How does that help? :S

OpenStudy (anonymous):

if you take out 6x you get: 6x*(-x+2)

OpenStudy (anonymous):

plz write the complete expression

OpenStudy (anonymous):

write down the complete exp...

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

\[6x^2=12x......6x=12...........x=2\]

OpenStudy (anonymous):

it's simple -6x²+12x=0 <=> -6x(1-2x)=0 <=> 1-2x=0 <=> x=1/2

OpenStudy (anonymous):

x= 1/2 that's it

OpenStudy (mayankdevnani):

\[-6x^2+12x\] \[6x(-x+2)\]

OpenStudy (mayankdevnani):

@Jon__Persson that's answer